OpenStudy (anonymous):

pre-calculus question x + 3^x < 4 from the graph or by testing you can see that x<1 but is there any algebrical solution to this?i mean with logs for example

OpenStudy (amistre64):

logs

OpenStudy (anonymous):

how?!

OpenStudy (anonymous):

there is no identity for log(x+3^x)

OpenStudy (amistre64):

thats what im thinking, have to iron out the thought now :)

OpenStudy (amistre64):

if we base each side \[e^{x+3^x}=e^4\] \[e^{x}e^{3^x}=e^4\] hmmm

OpenStudy (amistre64):

i got nothing at the moment. maybe a ratio of derivative ... but ive never grasped that concept

OpenStudy (anonymous):

its from a book that the derivatives are after that chapter and should be easy..so i suppose its from the graph or by testing and the solutions manual says just x<1

OpenStudy (amistre64):

setting up a linear approximation and taking the limit of it should get us there (just fancy trial and error really), but without a derivative to work with its just your usual trial and error

OpenStudy (anonymous):

yeah and the graphs even are at +3 chapters after that so i think its just trial and error...ty

OpenStudy (amistre64):

youre welcome